A 68 g sample of C6H5OH (Phenol) gas has a volume of 1.32 L at a temperature of 98°C.
i. Calculate the pressure using the Ideal Gas Law.
ii. Calculate the corrected pressure using the Van der Waals Equation.
atm×L2 L a=22.609 mol2 b=0.1177 mol
1
Expert's answer
2020-04-19T07:09:26-0400

1. nRT=PV

n=m/M=68/94=0.936

P=2187kPa=21.5 atm

2. (P+(n2a/V2)(V-nb)=(P+15,0)(1.32-0.11)=RT

P=25,2 atm